OpenStudy (anonymous):

x-2y=-4 3x+2y=12 solve by substitution

OpenStudy (blurbendy):

if you solve for x in the first equation you get: x = 2y -4 Now, substitute x = 2y -4 in for x in the bottom equation 3(2y-4) + 2y = 12 Solve for y Then put the answer you get for y back into the first equation, and solve for x

OpenStudy (anonymous):

how do you graph it ? it tells me to find the point where they intersect

OpenStudy (blurbendy):

whatever you get for x and y is the point (x,y) where the will intersect

OpenStudy (blurbendy):

they*

OpenStudy (blurbendy):

put both equations in y = mx +b form the first one would be -2y = -x -4 y = 1/2x + 2

OpenStudy (blurbendy):

now you know the slope for the first one (1/2)

OpenStudy (blurbendy):

start putting in points for x and y, so start with 1. put 1 in for x, then see what you get for y then you can plot your first point (x,y) keep doing that until your line is long enough. then do it for the second one. then see where they intersect